# Are We Learning the Right Features? A Framework for Evaluating DL-Based Software Vulnerability Detection Solutions

### v3

@@ V-B RQ2: Analysis of Feature Satisfaction Rate
For SFs that belong to Token-based sub-class, we observe that both LineVul and SySeVR produce around the same . For SFs that belong to graph-based sub-class, we observe that DeepWukong achieves a slightly higher (5 percentage points more in Node Set and 2 percentage points in Edge Set) than ReVeal, while DeepDFA achieves near perfect for both SFs.
@@ V-C RQ3: Analysis of FPP and FEP Satisfaction Rate
Graph-based | Token-based | ||||
VF | DeepWukong | ReVeal | DeepDFA | LineVul | SySeVR |
@@ VI Discussion and Implications
The evaluation of five detectors by VIPer yielded valuable insights, revealing the detectors’ respective strengths and weaknesses. In the case of DeepWukong, VIPer revealed that its lowest comes from VFs Double-Free and Use-After-Free which both belong to the sub-class Deallocated Buffer Use. Given that these vulnerabilities are primarily characterized by program control flow [25], VIPer’s findings suggest that DeepWukong struggles to comprehend control dependencies and their contribution to vulnerability. VIPer also shows that DeepWukong is also influenced by the SF Edge Set, which may be a key factor underlying its limited understanding of control dependencies.
For ReVeal, VIPer’s findings indicate it struggles to understand the impact of system APIs on vulnerabilities as it displays the lowest for VF Write API. ReVeal also shows a drop in recall for Node Set and Edge Set SFs. VIPer’s findings suggest that these SFs may impede ReVeal’s ability to recognize crucial dependencies related to system API calls.
@@
For LineVul, VIPer’s findings indicate that it achieves a but a . This disparity suggests that LineVul severely overfits when trained on SARD, excelling in false positive reduction but failing to generalize effectively. VIper also shows that the SF, Identifier Name impacts LineVul’s predictions by reducing its precision. This SF is one of the contributors to LineVul’s overfitting to SARD. Therefore, we suggest that LineVul should use some preprocessing technique on the tokens (e.g., token symbolizing) to avoid exposing itself to the SF Identifier Name. For SySeVR, VIPer shows that it gets lower compared to the three other detectors for most of the features. However, the lowest value is from the VF, Write API, suggesting that SySeVR’s token symbolization may inadvertently capture system APIs, hindering its ability to learn from these API names and understand their contribution to vulnerabilities. We suggest that SySeVR should update its token symbolization to ensure it only symbolizes user-defined functions.
@@ VII-1 External validity
This threats refers to the generalizability of the experiments and VIPer. We mitigated this threat by conducting a large-scale study including five recent and representative vulnerability detectors following token-based and graph-based approaches.555We additionally investigated the possibility of using a sixth vulnerability detector, DeepVD[34]. However, we were unable to reproduce DeepVD’s published results and did not receive a response from its authors when we asked for clarification. For this reason, we excluded DeepVD from our study. Although, the dataset employed in this study only contains C/C++ code samples, VIPer’s primary goal is to ensure accurate representation of vulnerabilities that should be learnt by the detectors, thus, it is language agnostic and dataset independent. Another threat may arise from our reliance on the SARD dataset when devising SFs. However, SARD is the largest available dataset that is widely used and contains many real-world security flaws. To further confirm that our results are not inadvertently impacted by SARD, we also examined how the graph-based SFs influence ReVeal by utilizing the FFmpeg+Qemu dataset. Our findings indicate that both SFs also affect ReVeal’s predictions on FFmpeg+Qemu, in a manner consistent with our SARD-based results. This supports the SFs’ broad applicability.
@@ VIII Related work
Understanding the features relevant to vulnerability prediction (i.e., VFs) is critical for improving the DL-based vulnerability detectors’ reliability and trustworthiness. Recent studies have, therefore, attempted to explore the VFs in DL-based vulnerability detectors. Risse et al. [18] studied the vulnerability detectors’ ability to learn fixes for vulnerable patches. Suneja et al. [35] probed the signal-awareness of models used for vulnerability detection by reducing the input source code to the minimum tokens required to retain the prediction. Another way researchers pursue explaining the predictions of DL-based models is by examining the unintended pattern in the dataset that the models might learn erroneously (i.e., SFs). On that note, recent studies have explored the influence of variable names [36, 17] and method names [37, 17] in the prediction of the models of code. Risse et al. [18] analyzed the behavior of models of code on logic-preserving transformations.

### v5

@@ V-B RQ2: Analysis of Feature Satisfaction Rate
For SFs that belong to the Token-based sub-class, both LineVul and SySeVR produce roughly the same . For SFs that belong to graph-based sub-class, DeepWukong achieves a slightly higher than ReVeal, while DeepDFA achieves near perfect for both SFs.
@@ V-C RQ3: Analysis of FPP and FEP Satisfaction Rate
| Graph-based | Token-based | |||||||||
| VF | DeepWukong | ReVeal | DeepDFA | LineVul | SySeVR | |||||
@@ VI Discussion and Implications
The evaluation of five detectors by VIPer yielded valuable insights, revealing the detectors’ respective strengths and weaknesses. In the case of DeepWukong, VIPer revealed that its lowest comes from VFs Double-Free and Use-After-Free, which both belong to the sub-class Deallocated Buffer Use. Given that these vulnerabilities are primarily characterized by program control flow [26], VIPer’s findings suggest that DeepWukong struggles to comprehend control dependencies and their contribution to vulnerability. VIPer shows that DeepWukong is also influenced by the SF Edge Set, which may be a key factor underlying its limited understanding of control dependencies.
VIPer’s findings indicate that ReVeal struggles to understand the impact of system APIs on vulnerabilities as it displays the lowest for VF Write API. ReVeal also shows a drop in recall for Node Set and Edge Set SFs. VIPer’s findings suggest that these SFs may impede ReVeal’s ability to recognize crucial dependencies related to system API calls.
@@
VIPer’s findings also indicate that LineVul achieves a but a . This disparity suggests that LineVul severely overfits when trained on SARD, excelling in false positive reduction but failing to generalize effectively. VIper also shows that the SF, Identifier Name impacts LineVul’s predictions by reducing its precision. This SF is one of the contributors to LineVul’s overfitting to SARD. Therefore, LineVul may benefit from using some preprocessing technique on the tokens (e.g., token symbolizing) to avoid exposing itself to the SF Identifier Name. For SySeVR, it gets lower compared to the four other detectors for most of the features. However, the lowest value is from the VF, Write API, suggesting that SySeVR’s token symbolization may inadvertently capture system APIs, hindering its ability to learn from these API names and understand their contribution to vulnerabilities. SySeVR may address this by modifying its token symbolization to include only user-defined functions.
@@ VII-1 External validity
This threats refers to the generalizability of the experiments and VIPer. We mitigated this threat by conducting a large-scale study including five recent and representative vulnerability detectors following token-based and graph-based approaches. 333We additionally investigated the possibility of using a sixth vulnerability detector, DeepVD[39]. However, we were unable to reproduce DeepVD’s published results and did not receive a response from its authors when we asked for clarification. For this reason, we excluded DeepVD from our study. Although, the dataset employed in this study only contains C/C++ code samples, VIPer’s primary goal is to ensure accurate representation of vulnerabilities that should be learnt by the detectors, thus, it is language agnostic and dataset independent. Another threat may arise from our reliance on the SARD dataset when devising SFs. However, SARD is the largest available dataset that is widely used and contains many real-world security flaws. To further confirm that our results are not inadvertently impacted by SARD, we also examined how the graph-based SFs influence ReVeal by utilizing the FFmpeg+Qemu dataset. Our findings indicate that both SFs also affect ReVeal’s predictions on FFmpeg+Qemu, in a manner consistent with our SARD-based results. This supports the SFs’ broad applicability.
@@ VIII Related work
Understanding the features relevant to vulnerability prediction (i.e., VFs) is critical for improving the DL-based vulnerability detectors’ reliability and trustworthiness. Recent studies have, therefore, attempted to explore the VFs in DL-based vulnerability detectors. Risse et al. [18] studied the vulnerability detectors’ ability to learn fixes for vulnerable patches. Suneja et al. [40] probed the signal-awareness of models used for vulnerability detection by reducing the input source code to the minimum tokens required to retain the prediction. Meanwhile, our approach for devising VFs is primarily focused on properties related to code dependencies. It is also the first to evaluate whether vulnerability detection techniques understand these code dependencies.
Another way researchers pursue explaining the predictions of DL-based models is by examining the unintended pattern in the dataset that the models might learn erroneously (i.e., SFs). On that note, recent studies have explored the influence of variable names [41, 17] and method names [42, 17] in the prediction of the models of code. Risse et al. [18] analyzed the behavior of models of code on logic-preserving transformations. Our work is the first to categorize SFs based on how they impact token-based and graph-based approaches, and to develop SFs based on graph properties of code. Additionally, this paper presents the dichotomy between VFs and SFs through an illustrative taxonomy as well as an evaluation framework that allows for the assessment of both VFs and SFs within the same platform.
@@ IX Conclusion and Future Work
Acknowledgment
This work is supported in part by the National Science Foundation under grant CCF-2106871.